# Activation Function Design Sustains Plasticity in Continual Learning

### v1

@@ Appendix F Expanding on Continual Reinforcement Learning Experiments
Activation Mean $\pm$ 95% CI Plasticity Score LR Optimal HP
ReLU 168.59  $\pm$ 222.92 0.0001 -
Leaky-ReLU -127.40  $\pm$ 73.09 1e-05 0.8
Sigmoid 993.61  $\pm$ 190.38 0.0001 -
Tanh 337.61  $\pm$ 20.01 1e-05 -
RReLU 435.00  $\pm$ 434.30 0.0001 Bounds: $[0.125,0.333]$
PReLU 684.88  $\pm$ 115.90 0.0001 Layer. $\alpha$=0.65
Swish (SiLU) 1120.28  $\pm$ 229.22 0.0001 0.01
GeLU 331.30  $\pm$ 126.07 0.0001 1.0
eLU -7.78  $\pm$ 194.69 1e-05 1.0
CeLU 115.78  $\pm$ 173.85 1e-05 1.0
SeLU 306.16  $\pm$ 4.27 1e-05 1.673
CReLU -69.09  $\pm$ 41.11 1e-05 -
Rational 603.39  $\pm$ 274.26 1e-05 C, 5, 4, Leaky-ReLU
Smooth-Leaky 1239.80  $\pm$ 212.96 0.0001 C:0.5, P:2.0, $\alpha$=0.1
Rand. Smooth-Leaky 1310.94  $\pm$ 98.37 0.0001 C:0.1, P:1.0. Bounds: $[0.01,0.1]$
@@
Table F1: Mean plasticity score across 5 seeds with 95% bootstrap confidence intervals (higher is better). Best mean per column is bolded. We also report optimal learning rate (LR) and optimal hyperparameters (HP) per activation function. In the case of bounded activations we provide the lower and upper bounds. PReLU provides granularity level at the number of parameters per activation layer. A dash (-), in the Optimal HP column, indicates that such activation function uses the unique or baseline parameter (e.g., ReLU only has slope $\alpha=0$).
@@
Previously, we defined $\mathrm{GAP}_{c,e}=R^{\text{train}}_{c,e}-R^{\text{test}}_{c,e}$, with $R$ the expected return measured at the end of cycle $c$. For each activation and environment, we summarize change across cycles as $\Delta(\mathrm{GAP}_{e})=\mathrm{GAP}_{3,e}-\mathrm{GAP}_{1,e}$ (abbreviated $\Delta$). Thus, $\Delta<0$ means the train–test gap shrinks across cycles (better transfer), whereas $\Delta>0$ means it widens (worse transfer).
We summarize across environments using both the median (robust) and the mean (sensitive to outliers). This distinction matters: for instance, GeLU exhibits an extremely large negative mean $\Delta$ driven by a single Humanoid-v5 outlier, whereas its median $\Delta$ is modest and positive (i.e., a typical slight widening), indicating that the outlier dominates the mean while the median reflects the typical behavior.
@@
Separately, we report a Plasticity Score that captures late-cycle functional performance on the training environments. We align runs to a common timeline, average rewards, and—within the last cycle—take the median of the mean returns over the final $15\%$ of each cycle. This metric answers: “can the agent still perform well after repeated shifts on the data it now collects?”—it is not a train–test gap.
The highest Plasticity Scores are obtained by Rand. Smooth-Leaky (1310.94), Smooth-Leaky (1239.80), Swish (1120.28), Sigmoid (993.61), and PReLU (684.88) (see Tab. F1 for full hyperparameter disclosure, we do not show Smooth-Leaky in favor of just showing the best of our custom activations). These activations reliably recover strong late-cycle returns across tasks. However, several also show positive $\Delta$ (e.g., Swish, Sigmoid, PReLU, Smooth-Leaky, Rand. Smooth-Leaky), indicating that while train performance rose, transfer to the perturbed tests did not keep pace (the gap widened). Conversely, Tanh and SeLU achieve negative median $\Delta$ (gaps narrow on typical environments) but only moderate Plasticity Scores ($\approx 337$ and $\approx 306$). Among ELU-family variants, eLU yields a small improvement in $\Delta$ with low train-side performance ($\approx 17$), whereas CeLU shows a larger gap reduction while showing modest competitiveness on Plasticity Score ($\approx 498.06$). ReLU and Leaky-ReLU produce low Plasticity Scores ($\approx 168.59$ / $\approx-80$), consistent with instability under repeated shifts, despite mixed signs in $\Delta$. Overall, a high Plasticity Score does not imply improved transfer: the score reflects adaptation on the training distribution, while $\Delta(\mathrm{GAP}_{\text{end}})$ reveals whether that adaptation carries to perturbed tests.
@@
We therefore present both viewpoints: (i) Plasticity Score for functional performance under non-stationarity, and (ii) Generalization Gap for evaluation of the adaptation carried to perturbed tests. We use the median $\Delta$ as the primary cross-environment summary and the mean $\Delta$ only to highlight outlier-driven effects. Full per-activation, per-environment results are in Tab. F2, and the cycle-by-cycle evolution is visualized in Fig. F1.
Activation HalfCheetah-v5 Humanoid-v5 Ant-v5 Hopper-v5 mean $\Delta$ median $\Delta$
ReLU 124.81 -1249177.00 183.00 -287.15 -312289.00 -81.17
Leaky-ReLU 65.64 -13959.34 -56.82 -0.48 -3487.75 -28.65
Sigmoid 1521.45 18.92 276.48 152.14 492.25 214.31
Tanh -782.69 48.28 -388.85 91.95 -257.83 -170.29
RReLU 527.09 -19274.62 -26.83 21.89 -4688.12 -2.47
PReLU 839.60 128132.00 94.17 -22.35 32260.85 466.88
Swish (SiLU) 627.24 -127118.70 1035.22 20.88 -31358.83 324.06
GeLU 317.71 -9766168.00 258.21 -32.29 -2441406.00 112.96
eLU 103.83 -80.15 -207.62 14.92 -42.25 -32.61
CeLU -341.13 -49.31 -281.51 3.56 -167.10 -165.41
SeLU -839.65 15.91 -339.88 51.90 -277.93 -161.99
CReLU 367.50 -65.10 -5.26 1.65 74.70 -1.81
Rational 550.81 9852.05 270.39 54.72 2681.99 410.60
Smooth-Leaky 621.48 -666714.76 -210.37 -8.04 -166577.9 -109.20
Rand. Smooth-Leaky 103.80 -1236645.00 563.09 19.22 -308989.80 61.51
@@
Table F2: First-to-last cycle $\Delta$ of the Generalization Gap per environment and activation function (rounded to 2 decimals). Lower is better (more transferability / higher plasticity). Best (lowest) per column is bolded.
@@ F.1 Interpreting Negative Generalization Gaps in Continual RL
A strongly negative generalization gap at the end of training ($\mathrm{GAP}_{c,e}<0$) is not paradoxical under non-stationary streams (e.g., randomized MuJoCo friction Abbas et al. (2023)). As the training environment keeps shifting, the agent can lose plasticity—it struggles to re-fit the current regime—so $R^{\text{train}}_{c,e}$ is depressed. Yet the policy may retain robust, regime-invariant skills that carry to perturbed test conditions, where evaluation is noise-free and does not suffer on-policy update instability. Consequently $R^{\text{test}}_{c,e}$ can exceed $R^{\text{train}}_{c,e}$, yielding a negative $\mathrm{GAP}_{\text{c,e}}$. Our summary $\Delta(\mathrm{GAP}_{e})=\mathrm{GAP}_{3,e}-\mathrm{GAP}_{1,e}$ becomes highly negative when transferability improves over the training cycle even as within-cycle adaptation degrades—an acceptable and informative outcome in this setting.
Figure F1: The heatmap reports end-of-cycle $\mathrm{GAP}_{c,e}$ per activation (rows) and cycle (columns). Colors are centered at 0 (green = negative values, test >> train; red = positive values, train >> test). Values are computed on a held-out friction variant of the training environment (Abbas et al., 2023). See Tab. F2 for the across-cycle summary $\Delta(\mathrm{GAP}_{e})$. Together, these views reveal when apparent trainability gains translate (or fail to translate) into generalization.
@@ G.2 Randomized-Slope SELU (RSELU)
During inference, the randomization is removed to ensure deterministic output, and the random variable $r$ is fixed to the mean of its distribution, $(l+u)/2$.
@@
A crucial feature of this design is that the bounds $l$ and $u$ are chosen to be symmetric around the original SELU alpha parameter ($\approx 1.6732$). This ensures that the self-normalizing property of SELU is preserved "in expectation" throughout training, providing a stable foundation while the slope randomization encourages robust learning and plasticity.

### v2

@@ Appendix F Expanding on Continual Reinforcement Learning Experiments
Activation IQM $\pm$ 95% CI Plasticity Score LR Optimal HP
ReLU 0.1593  $\pm$ 0.043 0.0001 -
Leaky-ReLU 0.1580  $\pm$ 0.115 1e-05 0.8
Sigmoid 0.3301  $\pm$ 0.059 0.0001 -
Tanh 0.2121  $\pm$ 0.028 1e-05 -
RReLU 0.2255  $\pm$ 0.059 0.0001 Bounds: $[0.125,0.333]$
PReLU 0.2695  $\pm$ 0.038 0.0001 Layer. $\alpha$=0.65
Swish (SiLU) 0.3130  $\pm$ 0.071 0.0001 0.01
GeLU 0.1658  $\pm$ 0.035 0.0001 1.0
eLU 0.1476  $\pm$ 0.107 1e-05 1.0
CeLU 0.1431  $\pm$ 0.070 1e-05 1.0
SeLU 0.2182  $\pm$ 0.032 1e-05 1.673
CReLU 0.1200  $\pm$ 0.010 1e-05 -
Rational 0.2217  $\pm$ 0.018 1e-05 C, 5, 4, Leaky-ReLU
SwiGLU 0.0781  $\pm$ 0.008 0.0001 -
Deep Fourier 0.1766  $\pm$ 0.008 1e-05 -
Smooth-Leaky 0.3283  $\pm$ 0.037 0.0001 C:0.5, P:2.0, $\alpha$=0.1
Rand. Smooth-Leaky 0.3853  $\pm$ 0.038 0.0001 C:0.1, P:1.0. Bounds: $[0.01,0.02]$
@@
Table F1: IQM plasticity score across 5 seeds with 95% bootstrap confidence intervals (higher is better). Best IQM per column is bolded. We also report optimal learning rate (LR) and optimal hyperparameters (HP) per activation function. In the case of bounded activations we provide the lower and upper bounds. PReLU provides granularity level at the number of parameters per activation layer. A dash (-), in the Optimal HP column, indicates that such activation function uses the unique or baseline parameter (e.g., ReLU only has slope $\alpha=0$).
Robust Metrics for Non-Stationary RL.
@@
Previously, we defined the generalization gap as $\mathrm{GAP}_{c,e}=R^{\text{train}}_{c,e}-R^{\text{test}}_{c,e}$. For each activation and environment, we summarize the change across cycles as $\Delta(\mathrm{GAP}_{e})=\mathrm{GAP}_{3,e}-\mathrm{GAP}_{1,e}$. A value of $\Delta<0$ implies the gap shrinks (improved transfer), while $\Delta>0$ implies it widens (worse transfer).
To provide a rigorous summary across environments with vastly different reward scales (e.g., Humanoid vs. Hopper), we depart from simple means or medians, which can be sensitive to outliers or mask the magnitude of failures. Instead, we adopt the Interquartile Mean (IQM) (Agarwal et al., 2021) for all cross-environment aggregations. Crucially, we address the issue of physics simulation instabilities (where plasticity loss leads to rewards $\ll-10^{6}$). We apply a stability filter: any run resulting in a physics explosion is treated as a functional failure. For the Generalization Gap, this means assigning a gap of $0.0$ (as no meaningful performance difference exists between two failures), preventing numerical artifacts from skewing the aggregate $\Delta$. We report the average reward per environment for all activation functions in Table F2.
Activation HalfCheetah-v5 Humanoid-v5 Ant-v5 Hopper-v5
ReLU $730.7\pm 787.7$ $1.0\times 10^{6}\pm 1.7\times 10^{6}$ $1041.7\pm 313.6$ $6004.2\pm 1.2\times 10^{4}$
Leaky-ReLU $2.5\times 10^{5}\pm 4.3\times 10^{5}$ $2.0\times 10^{5}\pm 1.3\times 10^{5}$ $4.5\times 10^{4}\pm 9.1\times 10^{4}$ $107.0\pm 62.8$
Sigmoid $2113.2\pm 811.5$ $\mathbf{342.2\pm 46.4}$ $1769.3\pm 243.6$ $\mathbf{336.2\pm 97.9}$
Tanh $1896.2\pm 578.8$ $194.9\pm 36.0$ $234.9\pm 126.7$ $109.7\pm 47.5$
RReLU $1260.8\pm 1259.2$ $9.3\times 10^{4}\pm 6.5\times 10^{4}$ $1151.3\pm 200.7$ $159.9\pm 125.1$
PReLU $2123.3\pm 1019.1$ $5.5\times 10^{5}\pm 4.5\times 10^{5}$ $1525.3\pm 425.5$ $146.2\pm 13.8$
Swish (SiLU) $2161.9\pm 806.8$ $7.6\times 10^{5}\pm 4.2\times 10^{5}$ $2135.2\pm 614.1$ $106.8\pm 87.1$
GeLU $850.8\pm 457.7$ $1.5\times 10^{7}\pm 2.6\times 10^{7}$ $766.1\pm 144.8$ $31.9\pm 41.3$
eLU $274.9\pm 2997.3$ $8.3\times 10^{5}\pm 8.6\times 10^{5}$ $4.1\times 10^{4}\pm 7.4\times 10^{4}$ $116.5\pm 102.7$
CeLU $521.3\pm 897.6$ $194.6\pm 96.4$ $42.2\pm 248.5$ $49.1\pm 75.4$
SeLU $1437.2\pm 618.1$ $270.0\pm 12.7$ $342.3\pm 17.4$ $174.0\pm 38.4$
CReLU $472.6\pm 684.6$ $1.0\times 10^{5}\pm 8.2\times 10^{4}$ $823.7\pm 165.7$ $12.1\pm 4.6$
Rational $1526.4\pm 317.9$ $6.1\times 10^{4}\pm 3.2\times 10^{4}$ $1083.3\pm 368.6$ $174.9\pm 102.3$
SwiGLU $4.6\times 10^{10}\pm 6.0\times 10^{10}$ $2.1\times 10^{30}\pm 3.3\times 10^{30}$ $05.8\pm 106.8 11$$18.0\pm 27.4$
Deep Fourier $592.5\pm 125.7$ $176.3\pm 41.5$ $327.7\pm 15.2$ $185.9\pm 47.0$
Smooth-Leaky $622.4\pm 536.5 2.2$$\times 10^{6}\pm 2.2\times 10^{6} 22$$202.9\pm 533.6$ $173.8\pm 39.6$
Rand. Smooth-Leaky $\mathbf{3221.5\pm 922.6}$ $6.0\times 10^{5}\pm 2.6\times 10^{5}$ $\mathbf{2791.2\pm 320.2}$ $187.2\pm 19.4$
Table F2: Average reward per environment (Mean $\pm$ Std Dev) over 5 seeds. Higher is better. Best performance per environment is bolded. Large negative values (e.g., for SwiGLU or Humanoid) indicate potential issues.
Plasticity Score Analysis.
@@
Separately, we report a Plasticity Score that captures late-cycle functional performance on the training environments. This is distinct from the generalization gap; it answers: “can the agent still perform well after repeated shifts on the data it now collects?” To make this metric comparable across tasks, we normalize returns to $[0,1]$ using robust global bounds derived from the entire experimental suite. We apply a stability floor to clipped rewards (treating physics failures as $0.0$) and report the IQM across seeds and environments.
Under this rigorous metric, the highest Plasticity Scores are obtained by Rand. Smooth-Leaky ($0.388$) and Sigmoid ($0.340$), followed by Smooth-Leaky ($0.330$) and Swish ($0.315$) (see Tab. F1 for full details). This ranking highlights a critical trade-off between peak plasticity and safety:
  • Rand. Smooth-Leaky achieves the highest aggregate score by dominating in solvable locomotion tasks (Ant, HalfCheetah), demonstrating that smooth, randomized non-linearities facilitate superior gradient flow and rapid adaptation. However, it lacks an upper bound, which leads to divergence in the volatile Humanoid environment.
  • Sigmoid achieves the second-best score via stability. Its bounded nature $(0,1)$ prevents physics explosions in Humanoid, securing a baseline of performance where others fail. However, this saturation limits its peak learning capacity in simpler environments, resulting in a lower total score than the randomized variant.
Regarding transfer, high plasticity often correlates with a widening gap. Sigmoid and Swish show positive IQM $\Delta$ (gaps widen), suggesting that their adaptation is somewhat specific to the current stationary distribution. Rand. Smooth-Leaky, while failing in Humanoid, actually exhibits the lowest IQM $\Delta$ among high-performers in the environments where it remains stable, suggesting its randomized landscape encourages more generalizable solutions. Conversely, traditional activations like ReLU and Leaky-ReLU produce low Plasticity Scores, consistent with their known instability under repeated shifts.
Table F1 provides the raw absolute rewards for all activations. This data is essential for contextualizing the Generalization Gap (Tab. 4); a low gap should only be considered a "success" if the corresponding absolute reward in Tab. F1 indicates the agent has actually solved the task.
@@
We therefore present both viewpoints: (i) Plasticity Score (IQM) for functional performance under non-stationarity, and (ii) Generalization Gap (IQM $\Delta$) for evaluation of the adaptation carried to perturbed tests. Full per-activation, per-environment results are in Tab. F3, and the cycle-by-cycle evolution is visualized in Fig. F1.
Activation HalfCheetah-v5 Humanoid-v5 Ant-v5 Hopper-v5 IQM $\Delta$ Mean $\Delta$ Std $\Delta$
ReLU 124.81  $\pm$ 511.73 203.06  $\pm$ 158.47 183.00  $\pm$ 473.82 -287.15  $\pm$ 547.78 153.91 55.93 231.12
Leaky-ReLU 546.91  $\pm$ 1425.72 0.00  $\pm$ 0.00 -168.86  $\pm$ 711.59 -1.91  $\pm$ 30.97 -0.95 94.04 312.12
Sigmoid -288.53  $\pm$ 2576.85 18.92  $\pm$ 111.38 276.48  $\pm$ 1018.48 152.14  $\pm$ 129.02 85.53 39.75 242.81
Tanh -467.58  $\pm$ 1506.19 -2.87  $\pm$ 52.27 21.61  $\pm$ 213.56 -49.31  $\pm$ 117.34 -26.09 -124.54 230.58
RReLU 527.09  $\pm$ 1341.00 38.56  $\pm$ 1398.73 -26.83  $\pm$ 530.91 21.89  $\pm$ 69.33 30.22 140.18 259.43
PReLU 839.60  $\pm$ 596.03 -316.28  $\pm$ 2211.72 94.17  $\pm$ 660.28 -22.35  $\pm$ 74.09 35.91 148.78 491.86
Swish (SiLU) 533.55  $\pm$ 1314.41 0.00  $\pm$ 0.00 780.79  $\pm$ 730.11 13.39  $\pm$ 37.27 273.47 331.93 388.92
GeLU 317.71  $\pm$ 730.89 -9.92  $\pm$ 245.15 258.21  $\pm$ 395.62 -32.29  $\pm$ 40.17 124.15 133.43 180.32
eLU 237.41  $\pm$ 740.39 617.50  $\pm$ 989.90 -81.21  $\pm$ 162.09 -10.55  $\pm$ 90.69 113.43 190.79 315.58
CeLU -341.13  $\pm$ 459.03 -49.31  $\pm$ 74.86 -281.51  $\pm$ 277.68 3.56  $\pm$ 44.91 -165.41 -167.10 169.68
SeLU 837.97  $\pm$ 1628.38 15.91  $\pm$ 80.82 -339.88  $\pm$ 185.25 51.90  $\pm$ 52.95 33.90 141.47 496.86
CReLU -238.91  $\pm$ 1039.91 640.56  $\pm$ 1255.50 329.34  $\pm$ 299.75 1.61  $\pm$ 3.40 165.48 183.15 383.71
Rational 550.82  $\pm$ 919.91 862.09  $\pm$ 2004.43 270.39  $\pm$ 526.94 54.72  $\pm$ 107.53 410.60 434.51 350.01
SwiGLU 0.00  $\pm$ 0.00 0.00  $\pm$ 0.00 -326.59  $\pm$ 186.38 -16.59  $\pm$ 44.16 -8.30 -85.80 160.72
Deep Fourier -568.95  $\pm$ 551.62 0.34  $\pm$ 58.99 -477.68  $\pm$ 208.27 41.53  $\pm$ 50.34 -238.67 -251.19 316.87
Smooth-Leaky 847.49  $\pm$ 1611.57 0.00  $\pm$ 0.00 44.09  $\pm$ 1089.34 27.01  $\pm$ 72.58 35.55 229.65 412.30
Rand. Smooth-Leaky -49.38  $\pm$ 883.36 0.00  $\pm$ 0.00 -336.13  $\pm$ 971.75 -68.68  $\pm$ 96.31 -59.03 -113.55 151.18
@@
Table F3: First-to-last cycle $\Delta$ of the Generalization Gap per environment and activation function (rounded to 2 decimals). The $\pm$ values indicate the 95% confidence interval. The IQM, mean, and std $\Delta$ are calculated across the 4 environments. Lower is better (negative values indicate the generalization gap decreased, improving plasticity). Best (lowest/most negative) per column is bolded.
@@ F.1 Interpreting Negative Generalization Gaps in Continual RL
A strongly negative generalization gap at the end of training ($\mathrm{GAP}_{c,e}<0$) is not paradoxical under non-stationary streams (e.g., randomized MuJoCo friction Dohare et al. (2024)). As the training environment keeps shifting, the agent can lose plasticity—it struggles to re-fit the current regime—so $R^{\text{train}}_{c,e}$ is depressed. Yet the policy may retain robust, regime-invariant skills that carry to perturbed test conditions, where evaluation is noise-free and does not suffer on-policy update instability. Consequently $R^{\text{test}}_{c,e}$ can exceed $R^{\text{train}}_{c,e}$, yielding a negative $\mathrm{GAP}_{\text{c,e}}$. Our summary $\Delta(\mathrm{GAP}_{e})$ becomes highly negative when transferability improves over the training cycle even as within-cycle adaptation degrades—an acceptable and informative outcome in this setting.
Figure F1: The heatmap reports end-of-cycle $\mathrm{GAP}_{c,e}$ per activation (rows) and cycle (columns). Colors are centered at 0 (green = negative values, test >> train; red = positive values, train >> test). Values are computed on a held-out friction variant of the training environment. See Tab. F3 for the across-cycle summary $\Delta(\mathrm{GAP}_{e})$. Together, these views reveal when apparent trainability gains translate (or fail to translate) into generalization.
@@ G.1 Bounded PReLU (Bo-PReLU)
PReLU 82.62  $\pm$ 0.05 92.67  $\pm$ 0.23 96.86  $\pm$ 0.32 43.30  $\pm$ 0.61 82.37  $\pm$ 0.11
Bo-PReLU 84.23  $\pm$ 0.02 91.57  $\pm$ 0.11 98.41  $\pm$ 0.01 48.15  $\pm$ 1.20 85.72  $\pm$ 0.11
Rand. Smooth-Leaky 84.26  $\pm$ 0.02 93.33  $\pm$ 0.05 98.42  $\pm$ 0.01 57.01  $\pm$ 1.59 86.23  $\pm$ 0.13
Table G2: Total Average Online Task Accuracy (%) on Continual Supervised Benchmarks, averaged over 5 independent runs. Values are reported as mean $\pm$ standard deviation (SD). Statistical significance was determined using an independent two-sample Welch’s t-test (p << 0.05).
Rand. Smooth-Leaky is only statistically significant with respect to Bo-PReLU on Random Label MNIST, CIFAR 5+1 and Continual ImageNet making Bo-PReLU highly competitive and highlighting the strengths of creating activation functions following the principles studied on Section 3 and Section 4. Optimal Hyperparameters for Bo-PReLU are indicated in Table G4, while the rest are in Table E2
@@ G.2 Randomized-Slope SELU (RSELU)
During inference, the randomization is removed to ensure deterministic output, and the random variable $r$ is fixed to the mean of its distribution, $(l+u)/2$. A crucial feature of this design is that the bounds $l$ and $u$ are chosen to be symmetric around the original SELU alpha parameter ($\approx 1.6732$). This ensures that the self-normalizing property of SELU is preserved "in expectation" throughout training, providing a stable foundation while the slope randomization encourages robust learning and plasticity.
@@
SeLU 80.43  $\pm$ 0.16 79.95  $\pm$ 0.91 84.61  $\pm$ 2.07 49.07  $\pm$ 1.25 80.98  $\pm$ 0.49
R-SeLU 81.72  $\pm$ 0.04 65.60  $\pm$ 2.07 47.23  $\pm$ 0.65 39.90  $\pm$ 2.91 79.87  $\pm$ 0.13
Rand. Smooth-Leaky 84.26  $\pm$ 0.02 93.33  $\pm$ 0.05 98.42  $\pm$ 0.01 57.01  $\pm$ 1.59 86.23  $\pm$ 0.13
Table G3: Total Average Online Task Accuracy (%) on Continual Supervised Benchmarks, averaged over 5 independent runs. Values are reported as mean $\pm$ standard deviation (SD). Statistical significance was determined using an independent two-sample Welch’s t-test (p << 0.05).
R-SeLU [0.423, 2.923] $\|$ 0.001 [0.423, 2.923] $\|$ 0.0001 [0.423, 2.923] $\|$ 0.001 [1.6732, 1.6732] $\|$ 0.001 [1.6732,1.6732] $\|$ 0.001
Bo-PReLU layer, $\alpha$ = 0.75, [0.5, 1.0] $\|$ 0.001 neuron, $\alpha$ = 0.65, [0.3, 1.0] $\|$ 0.001 neuron, $\alpha$ = 0.75, [0.5, 1.0] $\|$ 0.001 layer, $\alpha$ = 0.75, [0.5, 1.0] $\|$ 0.001 neuron, $\alpha$ = 0.65, [0.3, 1.0] $\|$ 0.001
Table G4: Optimal Hyperparameters for extra custom activation function in each Continual Supervised Benchmark Problem. Represented as activation function shape parameter on the left side of the $\|$ symbol and the learning rate to the right. Bo-PReLU’s $\alpha$ indicates initial parameter value and $[l,u]$ bounds. R-SeLU only indicates bounds $[l,u]$.